OpenStudy (anonymous):

how do you find the prime factorization of 72

OpenStudy (anonymous):

Well what are two numbers you can think of which multiply to yield 72?

OpenStudy (anonymous):

keep factoring 72. easiest way is keep dividing by 2.

OpenStudy (anonymous):

72/2 = 36. 36/2 = 18. 18/2 = 9. 9/3 = 3 so your factors are: 2 2 2 3 3

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[72 = 8*9 = 2^3*3^2\]